# SAM Decoding: Speculative Decoding via Suffix Automaton

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionized natural language processing by unifying tasks into text generation, yet their large parameter sizes and autoregressive nature limit inference speed. SAM-Decoding addresses this by introducing a novel retrieval-based speculative decoding method that uses a suffix automaton for efficient and accurate draft generation. Unlike n-gram matching used by the existing method, SAM-Decoding finds the longest suffix match in generating text and text corpuss, achieving an average time complexity of $O(1)$ per generation step. SAM-Decoding constructs static and dynamic suffix automatons for the text corpus and input prompts, respectively, enabling fast and precise draft generation. Meanwhile, it is designed as an approach that can be combined with existing methods, allowing SAM-Decoding to adaptively select a draft generation strategy based on the matching length, thus increasing the inference speed of the LLM. When combined with Token Recycling, evaluations show SAM-Decoding outperforms existing model-free methods, achieving a speedup of 2.27×2.27$\times$2.27 × over autoregressive decoding on Spec-Bench. When combined with EAGLE2, it reaches a speedup of 2.49×2.49$\times$2.49 ×, surpassing all current approaches. Our code is available at ours repository.
large language model, speculative decoding
@@ 1 Introduction
The Transformer-based Large Language Model (Brown et al., 2020; Dubey et al., 2024; Yang et al., 2024) has become a foundation model in the field of natural language processing. Through the large language model (LLM), all natural language processing tasks are unified as text generation tasks, which greatly improves the convenience of solving natural language processing tasks. While LLM has shown good performance in various tasks and is easy to use, the large-scale parameters and text generation-based paradigm contained in LLM have brought challenges to the practical deployment, especially in terms of the inference speed of the LLM. LLM generates tokens step by step in an autoregressive manner. At each round of generation, LLM generates one new token based on the tokens that has been generated. In addition, for the GPUs used for deploy LLM, this process requires reading the huge number of parameters contained in the LLM from the High-Bandwise Memory (HBM) to the on-chip memory, but these parameters only operate on a small number of parameters corresponding to the new token, resulting in the inability to fully utilize the parallel capabilities of the modern GPU, thereby limiting the inference speed of the LLM.
To address the above problem, one effective strategy is to enhance the arithmetic intensity of the large language model (LLM) generation process, which involves reducing the number of generation rounds while increasing the number of tokens generated in each round. Building upon this concept, the method of speculative decoding has been introduced and has spurred a series of follow-up studies (Leviathan et al., 2023; Miao et al., 2024; Cai et al., 2024; Fu et al., 2024). Rather than generating a single token per step, these approaches employ a draft model to predict token sequence as draft. The predicted sequence are then verified against the LLM, and only the accurate tokens are adopted.
Most speculative decoding techniques employ various neural networks, such as smaller language models (LLMs) or embedding layers, as draft models, which we categorize as model-based methods. While these model-based methods effectively enhance processing speeds, they are not without limitations. Firstly, the application of model-based speculative decoding algorithms necessitates the training of draft models, and this training must often be domain-specific, adding a layer of complexity to their practical deployment. Secondly, these preliminary models compete with the main LLMs for GPU memory resources. Additionally, utilizing a transformer architecture as a preliminary model introduces extra key-value caching requirements, further complicating GPU memory management within LLM inference systems (Kwon et al., 2023; Zheng et al., 2024).
@@
In addition to model-based speculative decoding methods, there is an alternative approach known as retrieval-based methods. These methods involve retrieving potential subsequent tokens from previously generated content or a text database during each generation step. Unlike model-based approaches, retrieval-based methods do not necessitate the training of draft models or the use of GPUs for inference. They are especially advantageous in tasks like summarization and code rewriting.
After examining the current retrieval-based methods, we identify several limitations. Firstly, these approaches typically rely on a single retrieval source. For instance, PLD (Saxena, 2023) focuses solely on prompts, whereas REST (He et al., 2024) is limited to text databases. Secondly, the retrieval techniques employed are often inefficient; the reliance on n-gram matching restricts the maximum length of matches, which can result in the production of less optimal drafts. Lastly, while retrieval-based methods excel in tasks like summarization, they generally underperform compared to model-based methods in other areas, limiting their piratical use.
@@
To address the limitations of the retrieval-based approach mentioned earlier, this paper introduces SAM-Decoding, an innovative retrieval-based speculative decoding technique. This method boasts more efficient and precise retrieval capabilities compared to traditional retrieval-based techniques. Unlike conventional approaches that rely on n-gram matching, SAM-Decoding employs a suffix automaton (SAM) to solve the longest suffix match problem for the generating text within both the historical generated text and the text database. This process yields a more accurate draft compared to n-gram matching. Moreover, during the text generation phase, the suffix automaton effectively captures the relationships between adjacent suffixes, ensuring that the average time complexity for identifying the longest suffix match at each step remains at $O(1)$. This represents a significant improvement over the computational efficiency of n-gram matching. Meanwhile, the length of the longest suffix match serves as an indicator of the alignment between the generating text and retrieval sources. By exploiting this metric, our approach can integrate existing speculative decoding methods as an auxiliary and dynamically decide whether to generate a draft based on the matched content, or utilize an auxiliary decoding method. This adaptability leads to a notable increase in the overall speed of the text generation process.
@@
Specifically, SAM-Decoding first constructs a static suffix automaton for the existing text corpus. For each request, SAM-Decoding then builds a dynamic suffix automaton based on the input prompt. The suffix automaton contains a series of nodes, each of which corresponds to several substrings in the retrieval sources. We record the earliest corresponding position of each node in the retrieval sources. This information facilitates the rapid generation of drafts by directly mapping nodes in the automaton to indices in the generating text as well as the text corpus. We maintain the corresponding nodes of the generating text in both the static and dynamic automaton. Therefore, in each round of generation, we can directly retrieve the draft from the retrieval sources according to the position recorded in the node, and filter the draft according to the matching length. After generation, as new tokens are generated, for the static automaton, we transition nodes based on these tokens. For the dynamic automaton, we first expand the structure to accommodate the new tokens before transitioning nodes accordingly.
@@
Extensive evaluations across a range of tasks indicate that the SAM-Decoding can achieve results that are competitive with existing state-of-the-art (SOTA) approaches in all tasks. Particularly in tasks where retrieval-based methods are applicable, SAM-Decoding outperforms all existing methods. On Spec-Bench, the latest benchmark for speculative decoding, SAM-Decoding combined with Token Recycling (Luo et al., 2024) achieves a speedup of 2.26 times compared to autoregressive decoding, outperforming all model-free baselines. Additionally, when SAM-Decoding is combined with EAGLE2 (Li et al., 2024b), the SOTA model-based speculative decoding method, it achieves a speedup of 2.49 times compared to autoregressive decoding, outperforming all existing model-based and model-free methods.
@@ 2 Background
In this section, we introduce background information relevant to this work.
@@ 2.1 Suffix Automaton
Suffix Automaton is an efficient data structure for representing the substring index of a given string, which allows the storage, processing, and retrieval of compressed information about all its substrings. The time complexity of constructing a suffix automaton is $O(L)$, where $L$ is the length of the string and it can be constructed incrementally.
A suffix automaton contains a series of nodes and two types of state transfer edges, next and link, where a node in the automaton represents a state and corresponds to a number of substrings that have the same ending position in the string. Meanwhile, two types of transfer edges are used to add a new token to the current state (next) and to take the suffix of the substring corresponding to the current state (link), thus transferring to a new state.
@@ 2.2 Speculative Decoding
LLM generates tokens in an autoregressive manner. Assuming that the input to the model is $x=(x_{1},x_{2},\ldots,x_{t})$, an LLM generates a new token at each generation step as follows:
$x_{t+1}=\text{Sample}(\text{LLM}(x)[t]).$ (1)
Speculative decoding improves the generation speed of LLMs by increasing the number of tokens generated per step and reducing the total number of generation steps. At each generation step, the speculative decoding method first generates a draft sequence $x_{\text{draft}}=(x_{t+1},x_{t+2},\ldots,x_{t+n})$ containing multiple tokens based on the draft model and the input $x$. The drafts are then verified by the LLM, and the correct tokens $x_{\text{accept}}=(x_{t+1},x_{t+2},\ldots,x_{t+m})$ are accepted:
$x_{draft}$ $=\text{Draft}(x),$
$x_{accept}$ $=\text{Verify}(\text{LLM}(x,x_{draft})[t:t+m]).$
By checking whether each token in the draft is the predicted result of the previous tokens in the LLM, we can filter out the accepted tokens. Specifically, we find the maximum number of consecutive tokens $m_{\text{accept}}$ such that:
$\text{Sample}(\text{LLM}(x,x_{draft})[t+i-1])=x_{draft}[i]$
$\quad i\in\{1,2,...,m_{accept}\}.$
In the above, we assume that the draft is a sequence of tokens. Recent works on speculative decoding have also proposed the use of tree-structured drafts. Specifically, these works introduce a tree mask in the attention module to compute tree-structured drafts without changing the existing framework. Using tree-structured drafts allows the LLM to validate multiple possible token sequences simultaneously, thereby increasing the acceptance rate of the drafts.
@@ 3 SAM-Decoding
In this section, we introduce our proposed method, SAM-Decoding. SAM-Decoding is a retrieval-based speculative decoding method designed to address three key limitations found in existing retrieval-based speculative methods: 1. The use of less comprehensive retrieval sources. 2. The employment of less efficient retrieval methods and restrictions on matching lengths. 3. Subpar performance outside specialized domains.
To tackle the first two limitations, SAM-Decoding leverages suffix automaton, which significantly enhance the comprehensiveness and efficiency of the retrieval process while allowing for more flexible matching lengths. Following this, we detail how SAM-Decoding can be integrated with both model-free and model-based methods. By utilizing the precise matching information provided by the suffix automaton, our method not only overcomes the third limitation but also ensures effective performance improvements across a wide range of tasks.
@@ 3.1 Suffix Automaton Construction
A suffix automaton can be constructed in linear time using Blumer’s algorithm (Blumer et al., 1984). Since the suffix automaton is designed for a single reference string, when dealing with multiple strings in the text corpus, we concatenate them into a single long string using special symbols, such as an End-of-Sentence (EOS) token. We then construct a static suffix automaton for this concatenated string.
At the same time, we also made several modifications to the suffix automaton to better suit draft generation. Specifically, at each generation step, the generating text corresponds to a node in the automaton, representing the longest suffix match. We need to identify the matching position in either the text corpus or the generating text. Based on this position, we can use the subsequent tokens as potential drafts. To achieve this, at every node of the suffix automaton, we record the earliest position of all substrings corresponding to that node in the reference string. We define this position as min_endpos. By using the min_endpos, we can directly map node to the position in the reference string. The construction process of the suffix automaton is detailed in Algorithm 3 in Appendix A.1.
@@ 3.3 Update of Suffix Automaton
where the process of expanding the suffix automaton is detailed in Algorithm 3 in Appendix A.1.
@@ 4 Experiments
Lookahead 1.29×1.29$\times$1.29 × 1.03×1.03$\times$1.03 × 1.22×1.$2$ 1.07×1.07$\times$1.07 × 1.41×1.41$\times$1.41 × 1.13×1.13$\times$1.13 × $1.62$ 44.37 1.19×1.19$\times$1.19 ×
@@
Vicuna-7B SAM-Decoding[T] 2.48$\times$ 1.73$\times$ 2.86$\times$ 1.98$\times$ 2.44$\times$ 2.14$\times$ 3.03 85.73 2.27$\times$
@@
Models and tasks. We conducted experiments on Vicuna-7B-v1.3 (Zheng et al., 2023). We evaluated SAM-Decoding on Spec-Bench (Xia et al., 2024), HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and HARGID (Kamalloo et al., 2023). Spec-Bench is a comprehensive benchmark designed for assessing Speculative Decoding methods across diverse scenarios. It is based on six commonly used datasets, MT-Bench (Zheng et al., 2023), WMT14 DE-EN, CNN/Daily Mail (Nallapati et al., 2016), Natural Question (Kwiatkowski et al., 2019), GSM8K (Cobbe et al., 2021), and DPR (Karpukhin et al., 2020), including six aspects: Multi-turn Conversation, Translation, Summarization, Question Answering, Mathmatical Reasoning, and Retrieval-augmented Generation. In addition, MBPP, HumanEval, and HARGID are used to evaluate the speed of decoding methods in code generation tasks and contextual question answering tasks, respectively.
Baselines. We considered the following baseline methods, including the model-based method EAGLE (Li et al., 2024a, b), the model-free methods Token Recycling (Luo et al., 2024) and Lookahead Decoding (Fu et al., 2024), and the retrieval-based methods PLD (Saxena, 2023) and REST (He et al., 2024).
Metrics. We evaluated speculative decoding methods from the following aspects (Li et al., 2024a)
@@
Experiment setup. We conducted experiments on a server equipped with a 20-core CPU and a single NVIDIA RTX A6000 GPU (48GB). The experiments were implemented using PyTorch 2.3.0 and CUDA 12.1. For the models, we used the float16 data type and applied greedy decoding with a batch size of 1. Regarding hyperparameters, unless otherwise specified, $l_{\rm bias}$ and $l_{\rm threshold}$ were set to 5, and the size of the draft generated by the automaton was set to 40 tokens. For the auxiliary speculative decoding algorithms, we used the default configurations as described in their respective original papers. One exception was made for Token Recycling: instead of the draft tree structure used in the original paper, we employed a draft tree with a depth of 6 and containing 60 nodes.
@@
Experiment Results. Experiment results on Spec-Bench are shown in Table 1. In the Spec-Bench, which consists of six tasks, multi-turn conversation, summarization, and retrieval-augmented generation were identified as particularly amenable to retrieval techniques. The results indicate that integrating SAM-Decoding into these tasks led to notable speed improvements. Specifically, for Token Recycling, the speedup ratio for the three tasks rose from 1.92×1.92$\times$1.92 ×, 1.96×1.96$\times$1.96 ×, and 1.68×1.68$\times$1.68 × to 2.48×2.48$\times$2.48 ×, 2.86×2.86$\times$2.86 ×, and 2.14×2.14$\times$2.14 ×, respectively. With EAGLE, the speedup ratio increased from 2.63×2.63$\times$2.63 ×, 2.28×2.28$\times$2.28 ×, and 1.95×1.95$\times$1.95 × to 2.78×2.78$\times$2.78 ×, 2.65×2.65$\times$2.65 ×, and 2.10×2.10$\times$2.10 ×, respectively. For EAGLE2, the speedup factors went up from 2.87×2.87$\times$2.87 ×, 2.33×2.$3$, and 2.03×2.03$\times$2.03 × to 3.02×3.02$\times$3.02 ×, 2.76×2.76$\times$2.76 ×, and 2.23×2.23$\times$2.23 ×, respectively. However, for the remaining tasks, the drafts produced by the retrieval methods did not significantly aid speculative decoding and instead introduced extra computational costs, leading to slight decreases in performance. Despite this, when SAM-Decoding was combined with these existing methods, there was an overall increase in speedup ratios: from 2.83×2.83$\times$2.83 × to 2.98×2.98$\times$2.98 × for Token Recycling, from 3.57×3.57$\times$3.57 × to 3.77×3.$7$ for EAGLE, and from 2.38×2.38$\times$2.38 × to 2.49×2.49$\times$2.49 × for EAGLE2.
It is also noteworthy that in the summarization task, the model-free approach (SAM-Decoding[T]) outperformed the model-based method, achieving a speedup ratio of 2.86×2.86$\times$2.86 ×. This finding underscores the potential benefits of integrating SAM-Decoding with different auxiliary strategies, especially in tasks where retrieval can effectively support the decoding process.
Table 2 presents the experimental outcomes on the HumanEval, MBPP, and HAGRID datasets. For the coding datasets HumanEval and MBPP, we constructed static suffix automata based on the python-code-instructions-18k dataset. For the contextual question answering dataset HAGRID, we built a static suffix automaton using its own training set.
@@
   w/o Token Recycling 1.89 59.26 1.57×1.57$\times$1.57 ×

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized natural language processing by unifying tasks into text generation, yet their large parameter sizes and autoregressive nature limit inference speed. SAM-Decoding addresses this by introducing a novel retrieval-based speculative decoding method that uses a suffix automaton for efficient and accurate draft generation. Unlike n-gram matching used by the existing method, SAM-Decoding finds the exact longest suffix match in generating text and text corpuss, achieving an average time complexity of $O(1)$ per generation step. Meanwhile, based on the exact match length, it is designed as an approach that can be combined with existing methods, allowing SAM-Decoding to adaptively select a draft generation strategy based on the matching length, thus increasing the inference speed of the LLM. When combined with Token Recycling, evaluations show SAM-Decoding outperforms existing model-free methods, achieving a speedup of 2.27×2.27$\times$2.27 × over autoregressive decoding on Spec-Bench. When combined with EAGLE2, it reaches a speedup of 2.49×2.49$\times$2.49 ×, surpassing all current approaches. Our code is available at ours repository.
SAM Decoding: Speculative Decoding via Suffix Automaton
Yuxuan Hu1,2, Ke Wang1,2, Xiaokang Zhang1,2, Fanjin Zhang4 Cuiping Li1,3Hong Chen1,3Jing Zhang1,3thanks: Corresponding author. 1School of Information, Renmin University of China, Beijing, China 2Key Laboratory of Data Engineering and Knowledge Engineering, Beijing, China 3Engineering Research Center of Database and Business Intelligence, Beijing, China 4Knowledge Engineering Group, Tsinghua University, Beijing, China
@@ 1 Introduction
The Transformer-based Large Language Model Brown et al. (2020); Dubey et al. (2024); Yang et al. (2024) has become a foundation model in the field of natural language processing. Through the large language model (LLM), all natural language processing tasks are unified as text generation tasks, which greatly improves the convenience of solving natural language processing tasks. While LLM has shown good performance in various tasks and is easy to use, the large-scale parameters and text generation-based paradigm contained in LLM have brought challenges to the practical deployment, especially in terms of the inference speed of the LLM. LLM generates tokens step by step in an autoregressive manner. At each round of generation, LLM generates one new token based on the tokens that has been generated. In addition, for the GPUs used for deploy LLM, this process requires reading the huge number of parameters contained in the LLM from the High-Bandwise Memory (HBM) to the on-chip memory, but these parameters only operate on a small number of parameters corresponding to the new token, resulting in the inability to fully utilize the parallel capabilities of the modern GPU, thereby limiting the inference speed of the LLM.
To address the above problem, one effective strategy is to enhance the arithmetic intensity of the large language model (LLM) generation process, which involves reducing the number of generation rounds while increasing the number of tokens generated in each round. Building upon this concept, the method of speculative decoding has been introduced and has spurred a series of follow-up studies Leviathan et al. (2023); Miao et al. (2024); Cai et al. (2024); Fu et al. (2024). Rather than generating a single token per step, these approaches employ a draft model to predict token sequence as draft. The predicted sequence are then verified against the LLM, and only the accurate tokens are adopted.
Most speculative decoding techniques employ various neural networks, such as smaller language models (LLMs) or embedding layers, as draft models, which we categorize as model-based methods. Although model-based speculative methods perform well on all kinds of tasks, when we consider some special tasks such as summarization, rewriting, there is a simpler type of speculative decoding methods that exhibit superior performance, which we call retrieval-based speculative decoding methods. These methods involve retrieving potential subsequent tokens from previously generated content or a text database during each generation round. Unlike model-based approaches, retrieval-based methods do not necessitate the training of draft models or the use of GPUs for inference.
@@
After examining the current retrieval-based methods, we identify several limitations. Firstly, these approaches typically rely on a single retrieval source. For instance, PLD Saxena (2023) focuses solely on prompts, whereas REST He et al. (2024) is limited to text corpus. Secondly, there are limitations in the retrieval techniques they adopt; PLD relies on n-gram matching, which is fast for context matching but cannot be extended to a wider range of textual repositories due to its poor theoretical complexity, while REST is based on suffix arrays, which has better complexity but cannot be applied to continually generating text. Lastly, not every step in the generation is suitable for the retrieval method; in a few generation steps, a large number of tokens can be accepted based on PLD and REST, but in the rest of the generation steps there is no acceleration.
@@
To address the limitations of the retrieval-based approach mentioned earlier, this paper introduces SAM-Decoding, an innovative retrieval-based speculative decoding technique. Unlike existing approaches that rely on n-gram matching, SAM-Decoding employs a suffix automaton (SAM) to solve the exact longest suffix match problem for the generating text within both the historical generated text and the text corpus. This process yields more accurate match positions and exact match lengths compared to n-gram matching. Moreover, during the text generation phase, the suffix automaton effectively captures the relationships between adjacent suffixes, ensuring that the average time complexity for identifying the longest suffix match at each step remains at $O(1)$. This represents a significant improvement over the theoretical computational efficiency of n-gram matching. Meanwhile, the length of the longest suffix match serves as an indicator of the alignment between the generating text and retrieval sources. By exploiting this metric, our approach can integrate existing speculative decoding methods as an auxiliary and dynamically decide whether to generate a draft based on the matched content, or utilize an auxiliary decoding method. This adaptability leads to a notable increase in the overall speed of the text generation process.
@@
Specifically, SAM-Decoding first constructs a static suffix automaton for the existing text corpus. For each request, SAM-Decoding then builds a dynamic suffix automaton based on the input prompt. The suffix automaton contains a series of nodes, each of which corresponds to several substrings in the context (text corpus, prompt and generating text). We record the earliest corresponding position of each node in the text. This information facilitates the generation of drafts by directly mapping nodes in the automaton to indices in the context. We also maintain the corresponding nodes of the generating text in both the static and dynamic automaton. Therefore, in each round of generation, we can directly retrieve the draft from the context according to the position recorded in the node, and filter the draft according to the matching length. After generation, as new tokens are generated, for the static automaton, we transition nodes of SAM based on new tokens, and for the dynamic automaton, we first expand the structure to accommodate the new tokens before transitioning nodes of SAM accordingly.
@@
Extensive evaluations across a range of tasks indicate that the SAM-Decoding can achieve results that are competitive with existing state-of-the-art (SOTA) approaches in all tasks. Particularly in tasks where retrieval-based methods are applicable, SAM-Decoding outperforms all existing methods. On Spec-Bench, the latest benchmark for speculative decoding, SAM-Decoding combined with Token Recycling Luo et al. (2024) achieves a speedup of 2.26 times compared to autoregressive decoding, outperforming all model-free baselines. Additionally, when SAM-Decoding is combined with EAGLE2 Li et al. (2024a), the SOTA model-based speculative decoding method, it achieves a speedup of 2.49 times compared to autoregressive decoding, outperforming all existing model-based and model-free methods.
@@ 2.1 Suffix Automaton
Suffix Automaton is an efficient data structure for representing the substring index of a given string, which allows fast substring retrieval. The time complexity of constructing a suffix automaton is $O(L)$, where $L$ is the length of the string and it can be constructed incrementally.
A suffix automaton contains a series of nodes and two types of state transfer edges, extension edges and suffix link edges. A node in the automaton represents a state and corresponds to all substrings that have the same ending position in the string. Meanwhile, extension edges are standard edges that represent a possible extension of the current substring by appending a new character, while suffix link edges create a path that allows the automaton to quickly jump to states representing shorter suffixes of the current substring.
@@ 2.2 Speculative Decoding
Given the model input $x=(x_{1},x_{2},\ldots,x_{t})$, an LLM generates a new token $x_{t+1}$ at each generation step autoregressively. The key idea of speculative decoding is to utilize a lightweight draft model to generate multiple candidate tokens quickly, i.e., $x_{\text{draft}}=(x_{t+1},x_{t+2},\ldots,x_{t+n})$, and then the target LLM simultaneously evaluates these candidates and accept those aligned with the output distribution of the LLM, i.e., $x_{\text{accept}}=(x_{t+1},x_{t+2},\ldots,x_{t+m})$.
In the above, we assume that the draft is a sequence of tokens. Recent works proposed to generate a candidate token tree via a tree mask in the attention module to make the target LLM simultaneously evaluate multiple branches of this token tree, thereby increasing the acceptance length of the draft model.
@@ 3 SAM-Decoding
In this section, we introduce our proposed method, SAM-Decoding. SAM-Decoding is a retrieval-based speculative decoding method designed to address three key limitations in existing retrieval-based speculative methods: (1) The use of insufficient retrieval sources. (2) The employment of inefficient retrieval methods and restrictions on $n$-gram matching lengths. (3) Subpar performance outside specialized domains (e.g., summarization and RAG tasks).
To tackle the first two limitations, SAM-Decoding leverages suffix automaton, which significantly enhances the coverage of retrieved corpus and the efficiency of the retrieval process while allowing for more flexible matching lengths. In what follows, we detail how SAM-Decoding can be integrated with both model-free and model-based methods. By utilizing the precise matching information provided by the suffix automaton, our method not only overcomes the third limitation but also ensures consistent performance improvements across a wide range of tasks.
@@ 3.1 Suffix Automaton Construction
A suffix automaton can be constructed in linear time using Blumer’s algorithm Blumer et al. (1984). Since the suffix automaton is designed for a single reference string, when dealing with multiple strings in the text corpus, we concatenate them into a single long string using special symbols, such as an End-of-Sentence (EOS) token. We then construct a static suffix automaton for this concatenated string.
At the same time, we also made several modifications to the suffix automaton to better suit draft generation. Specifically, at each generation step, the generating text corresponds to a node in the automaton, representing the longest suffix match. We need to identify the matching position in either the text corpus or the generating text. Based on this position, we can use the subsequent tokens as potential drafts. To achieve this, at every node of the suffix automaton, we record the earliest position of all substrings corresponding to that node in the reference string. We define this position as min_endpos. By using the min_endpos, we can directly map node to the position in the reference string. The construction process of the suffix automaton is detailed in Appendix A.1.
@@ 3.3 Update of Suffix Automaton
where the process of expanding the suffix automaton is detailed in Appendix A.1.
@@ 4 Experiments
Vicuna-7B Lookahead 1.29×1.29$\times$1.29 × 1.03×1.03$\times$1.03 × 1.22×1.$2$ 1.07×1.07$\times$1.07 × 1.41×1.41$\times$1.41 × 1.13×1.13$\times$1.13 × $1.62$ 44.37 1.19×1.19$\times$1.19 ×
@@
SAM-Decoding 1.75$\times$ 0.99$\times$ 2.54$\times$ 1.24$\times$ 1.74$\times$ 1.71$\times$ 1.89 62.26 1.65$\times$
@@
SAM-Decoding[T] 2.48$\times$ 1.73$\times$ 2.86$\times$ 1.98$\times$ 2.44$\times$ 2.14$\times$ 3.03 85.73 2.27$\times$
@@
Models and tasks. We conducted experiments on Vicuna-7B-v1.3 Zheng et al. (2023). We evaluated SAM-Decoding on Spec-Bench Xia et al. (2024), HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and HARGID Kamalloo et al. (2023). Spec-Bench is a comprehensive benchmark designed for assessing Speculative Decoding methods across diverse scenarios. It is based on six commonly used datasets, MT-Bench Zheng et al. (2023), WMT14 DE-EN, CNN/Daily Mail Nallapati et al. (2016), Natural Question Kwiatkowski et al. (2019), GSM8K Cobbe et al. (2021), and DPR Karpukhin et al. (2020), including six aspects: Multi-turn Conversation, Translation, Summarization, Question Answering, Mathmatical Reasoning, and Retrieval-augmented Generation. In addition, MBPP, HumanEval, and HARGID are used to evaluate the speed of decoding methods in code generation tasks and contextual question answering tasks, respectively.
Baselines. We considered the following baseline methods, including the model-based method EAGLE Li et al. (2024b, a), the model-free methods Token Recycling Luo et al. (2024) and Lookahead Decoding Fu et al. (2024), and the retrieval-based methods PLD Saxena (2023) and REST He et al. (2024).
Metrics. We evaluated speculative decoding methods from the following aspects Li et al. (2024b)
@@
Experiment setup. We conducted experiments on a server equipped with a 20-core CPU and a single NVIDIA RTX A6000 GPU (48GB). The experiments were implemented using PyTorch 2.3.0 and CUDA 12.1. For the models, we used the float16 data type and applied greedy decoding with a batch size of 1. Regarding hyperparameters, unless otherwise specified, $l_{\rm bias}$ and $l_{\rm threshold}$ were set to 5, and the size of the draft generated by the automaton was set to 40 tokens. For the auxiliary speculative decoding algorithms, we used the default configurations as described in their respective original papers. One exception was made for Token Recycling: instead of the draft tree structure used in the original paper, we employed a draft tree with a depth of 6 and containing 60 nodes. In addition, for the vicuna model, we use the vicuna template to convert prompt and output into the chat form, and for the Llama3 model, we use the template provided in the EAGLE2 repository.
@@
Experiment Results. Experiment results on Spec-Bench when using Vicuna-7B-v1.3 are shown in Table 1. In the Spec-Bench, which consists of six tasks, multiturn conversation, summarization, and retrieval-augmented generation were identified as particularly amenable to retrieval techniques. The results indicate that integrating SAM-Decoding into these tasks led to notable speed improvements. Specifically, for Token Recycling, the speedup ratio for the three tasks rose from 1.92×1.92$\times$1.92 ×, 1.96×1.96$\times$1.96 ×, and 1.68×1.68$\times$1.68 × to 2.48×2.48$\times$2.48 ×, 2.86×2.86$\times$2.86 ×, and 2.14×2.14$\times$2.14 ×, respectively. With EAGLE, the speedup ratio increased from 2.63×2.63$\times$2.63 ×, 2.28×2.28$\times$2.28 ×, and 1.95×1.95$\times$1.95 × to 2.78×2.78$\times$2.78 ×, 2.65×2.65$\times$2.65 ×, and 2.10×2.10$\times$2.10 ×, respectively. For EAGLE2, the speedup ratios went up from 2.87×2.87$\times$2.87 ×, 2.33×2.$3$, and 2.03×2.03$\times$2.03 × to 3.02×3.02$\times$3.02 ×, 2.76×2.76$\times$2.76 ×, and 2.23×2.23$\times$2.23 ×, respectively. However, for the remaining tasks, the drafts produced by the retrieval methods did not significantly aid speculative decoding and instead introduced extra computational costs, leading to slight decreases in performance. Despite this, when SAM-Decoding was combined with these existing methods, there was an overall increase in speedup ratios: from 2.83×2.83$\times$2.83 × to 2.98×2.98$\times$2.98 × for Token Recycling, from 3.57×3.57$\times$3.57 × to 3.77×3.$7$ for EAGLE, and from 2.38×2.38$\times$2.38 × to 2.49×2.49$\times$2.49 × for EAGLE2. It is noteworthy that in the summarization task, the model-free approach (SAM-Decoding[T]) outperformed the model-based method, achieving a speedup ratio of 2.86×2.86$\times$2.86 ×. Meanwhile, these results also underscores the potential benefits of integrating SAM-Decoding with different auxiliary strategies, especially in tasks where retrieval can effectively support the decoding process.
Table 2 presents the experimental results on the HumanEval, MBPP, and HAGRID datasets when using Vicuna-7B-v1.3. For the coding datasets HumanEval and MBPP, we constructed static suffix automata based on the python-code-instructions-18k dataset. For the contextual question answering dataset HAGRID, we built a static suffix automaton using its own training set.
@@
In addition to Vicuna-7B, we also conducted experiments on more models, and the experimental results are shown in Appendix B.
@@
   w/o Token Recycling 1.89 62.26 1.65×1.65$\times$1.65 ×
@@
For more ablation experiment results, please refer to Appendix C.